OpenStudy (anonymous):

\[\int\limits 3x \sqrt {x+2} ~dx--->x=u^2-2\]\[2\int\limits 3(u^2-2) \sqrt {u^2}\times u\]How to carry on?

OpenStudy (experimentx):

multiply and simplify first

OpenStudy (experimentx):

if you know how to integrate polynomial then rest is a piece of cake

OpenStudy (anonymous):

let u=x+2

OpenStudy (anonymous):

then du=dx

OpenStudy (anonymous):

then u-2=x

OpenStudy (anonymous):



OpenStudy (anonymous):

now distribute the u^1/2 into the parentheses

OpenStudy (anonymous):

\[I~have~so~far:->~ 2 \int\limits 3u^4-6u^2\]

OpenStudy (anonymous):

how is that possibe, dont you see what i did up there ^

OpenStudy (experimentx):